OpenStudy (anonymous):

Find the equation of a line in standard form when m=3;b=-2

OpenStudy (theeric):

Hi! Have you got a clue where to start? Any knowledge you can just throw out here about this?

OpenStudy (anonymous):

no but all I know is that m=3 and b=-2 and you need to find the line in standard form...

OpenStudy (theeric):

Alright, we'll go off of that. I'll tell you what the standard form of a line is (according to some of the first links on Google):\[Ax+By=C\]Do you know what the m and b represent?

OpenStudy (anonymous):

y-y1=m(x-x1) this is how it is supposed to me then once you are done Im pretty sure you just plug it in but just don know what to do with b.

OpenStudy (theeric):

What you wrote is the "point-slope" form, where you have a point (x1, y1) and the slope, m. For this problem, you are given b, which is the y-intercept. It is used in "slope-intercept" form. Yes, there are many many ways to write the same line... But you have m and b! Are you aware of slope-intercept form?\[y=mx+b\]

OpenStudy (theeric):

If you know slope-intercept form, that's good! You can use algebra to turn it into the standard form!

OpenStudy (anonymous):

yea I know I just don't know were in the world I plug in the b

OpenStudy (theeric):

Let me show you the goal:\[y=mx+b \rightarrow\rightarrow\rightarrow Ax+By=C\]

OpenStudy (theeric):

so you want, IN GENERAL (it'll be very helpful to keep your mind in "in general" mode for a moment), something times y, plus something times x, equaling something (no y or x)

OpenStudy (theeric):

One of those somethings CAN be 1! And so you might not see it, but it's there.

OpenStudy (anonymous):

magic

OpenStudy (theeric):

Example (and this is a hint, as well):\[y=1y=By \text{ where }B\text{ is }1\]

OpenStudy (theeric):

Yes!\[\rightarrow\rightarrow\rightarrow=magic!!!\]

OpenStudy (theeric):

You will do this magic :)

OpenStudy (anonymous):

\[y+2=3(x-0)\] so I got it first I pluged everything in and I want you to tell me if I do anything wrong but then \[y+2=3x\] the you subtract the y so it can go on the other side which it would look like this ANSWER:2=3x-y

OpenStudy (theeric):

Yep! That's one way to do it! Congrats! All you have to do to check is look to make sure you have the \[A,B,C\]in\[C=Ax-By\]which you do! :)

OpenStudy (anonymous):

MAGIC

OpenStudy (theeric):

\[C=Ax+By\] I mean

OpenStudy (theeric):

As long as you personally can identify A, B, and C, you've got this problem!

OpenStudy (anonymous):

ok thank you Eric!

OpenStudy (theeric):

You're very welcome! Thank yourself! I did hardly anything! :)